OpenStudy (anonymous):

What is the solution to the system? Line 1 x y –2 –4 –1 –1 Line 2 x y 2 3 3 1 ( , )

OpenStudy (jdoe0001):

why not graph them both? both are LINEar equations, and thus their graph is a LINE to graph a line, all you need is two points and you have two for each plot them both, where they meet, or "intersect", that's the solution to the system

OpenStudy (anonymous):

I need help finding the y-intercept

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... you really don't need it, just to graph them though

OpenStudy (anonymous):

You need the y-intercept to graph, right?

OpenStudy (jdoe0001):

nope to graph a line, all you need is two points 

OpenStudy (anonymous):

So, all I need is the slope?

OpenStudy (jdoe0001):

well... you don't.... why not use the points given, and plot both functions?

OpenStudy (jdoe0001):

you could get the equations, sure and make them slope-intercept form, if you wish and use that will yield the same though

OpenStudy (anonymous):

Is it (1,5)?

OpenStudy (dtan5457):

Do you still need help on this?

OpenStudy (dtan5457):

If you put this both to a linear equation (I got from point slope formula) you get y=3x+2 and y=-2x+7 set these two together 3x+2=-2x+7 x=1 plug that back in to any of the linear equations, and you get 5. so yeah, your right, it's (1,5)